OpenStudy (anonymous):

Need some serious help in solving literal equations I'm pretty ok when it comes to solving inequalities but now I have so many variables I'm not sure what to do...how do I solve for this stuff? Solve for c. ab + c = 5

OpenStudy (anonymous):

You can't solve that without another equation...

OpenStudy (anonymous):

what do you mean? why would you need another equation?

OpenStudy (asnaseer):

ok - first step here would be to subtract ab from both sides - what do you end up with?

OpenStudy (anonymous):

c - 5 - ab?

OpenStudy (anonymous):

no I meant c = 5 - ab

OpenStudy (asnaseer):

what happened to the equals sign?

OpenStudy (asnaseer):

yes :)

OpenStudy (asnaseer):

and there you have magically solved for c :)

OpenStudy (anonymous):

I dont see how I can figure out what to do to solve for them. :/

OpenStudy (asnaseer):

you just need to keep an eye out for which variable you are solving for. then treat everything else as just "bundles" of unknown quantities

OpenStudy (asnaseer):

e.g. try solving this one for c: xyz + 2c = c - pq

OpenStudy (asnaseer):

remember the main aim is to get the symbol of interest (c in this case) over to the left-hand-side, and everything else over to the right-hand-side

OpenStudy (anonymous):

so I have to have the variable I'm solving for on the left side?

OpenStudy (asnaseer):

yes :)

OpenStudy (asnaseer):

it is always best to do these one step at a time

OpenStudy (asnaseer):

so first what do you think you would need to do in order to move the xyz to the right-hand-side?

OpenStudy (anonymous):

there are two c's. Do I have to combine like terms, or do I have to keep it that way because both cs are on different sides of the equal sign?

OpenStudy (asnaseer):

just do one step at a time - so do the first step I asked above

OpenStudy (anonymous):

do I subtract xyz from both sides?

OpenStudy (asnaseer):

perfect! - and what do you end up with?

OpenStudy (anonymous):

making it 2c = c - pq - xyz

OpenStudy (asnaseer):

great! now next step would be to get that lone c from the right-hand-side over to the left

OpenStudy (anonymous):

mm...I'm stuck...should I divide 2c and c by something?

OpenStudy (asnaseer):

think about what you need to do in order to remove the lone c from the right-hand-side, then apply that same process to the left

OpenStudy (anonymous):

I would subtract c, but if I did thta, it would remove c entirely from the equation, wouldn't it?

OpenStudy (asnaseer):

that is correct - so subtract c from both sides - what do you end up with?

OpenStudy (anonymous):

...2 = pq?

OpenStudy (asnaseer):

what happened there?

OpenStudy (asnaseer):

we started with: 2c = c - pq - xyz we now need to subtract c from both sides

OpenStudy (anonymous):

2c = pq - xyz, right?

OpenStudy (asnaseer):

remember that 2c - c = c

OpenStudy (anonymous):

why is that? if I subtracted c from 2c that would make it 2, right?

OpenStudy (asnaseer):

2c = c - pq - xyz -c = -c - pq - xyz ============ c = -pq - xyz

OpenStudy (asnaseer):

think of it like this: what is 2 apples take away one apple?

OpenStudy (anonymous):

oh! I see! there are two cs and I am subtracting a c. I forgot that a variable counts as one digit

OpenStudy (asnaseer):

good - I hope it is becoming clearer for you now :)

OpenStudy (anonymous):

yes, thank you. :]

OpenStudy (asnaseer):

yw :)